A particle having a charge of 2.0 × 10^-8 C and a mass of 2.0 × 10^-10 g is projected with a speed of 2.0 × 10^3 m/s. in a region having a uniform magnetic field of 0.10 T. The velocity is perpendicular to the field. Find the radius of the circle formed by the particle and also the time period.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. q = 2.0 × 10–8 C B ⃗ = 0.10 T m = 2.0 * 10^-10 g = 2 * 10^-13 g υ = 2.0 * 10^3 m/’ R = m υ/qB = 2 * 10^-13 *2 * 10^3/2 * 10^-8 * 10^-1 = 0.2 m = 20 cm T = 2πm/qB = 2 * 3.14 *2 * 10^-13/2 * 10^-18 * 10^-1 = 6.28 * 10^-4 s